OpenStudy (anonymous):

Let f(x) = x2 – 7x + 3 and g(x) = x + 4. Find f(x) + g(x).

OpenStudy (nali):

just add them 

OpenStudy (anonymous):

g(-1) = (-1)^2 - 7(-1) - 9 g(-1) = -2+7-9 g(-1) = -1 now you plug -1 (which is g(-1)) into f(x) in place of x. f(-1) = -1 - 2 f(-1) = -3